OpenStudy (anonymous):

The perimeter of a square exceeds that of another by 100cm and the area of the larger square exceeds three times the area of the smaller square by 325 cm^2. Find the length of a side of each of the squares .. Help is much appreciated :)

OpenStudy (anonymous):

4x = 4y + 100 x^2 = 3y^2 + 325 solve these system of equations.... x being the side of the bigger square.

OpenStudy (anonymous):

simplify that first equation to x=y+25 and substitute into the second equation....

OpenStudy (anonymous):

It's still confusing me :X Sorry Bad in maths

OpenStudy (anonymous):

ok... substituting x=y+25 into the second equation, it becomes (y+25)^2 = 3y^2 + 325 y^2 + 50y + 625 = 3y^2 + 325 -2y^2 + 50y - 300 = 0 2y^2 - 50y + 300 = 0 y^2 - 25y + 150 = 0 (y - 30)(y + 5) = 0 ===> y = 30 can you take it from here?

OpenStudy (anonymous):

Yes thank you :)